OpenStudy (anonymous):

Need help!! 8th grade math.. pleasee if you answer you get a medal Write each equation in slope-intercept form. Make a sketch of the graph. Include the x- and y-intercepts. 49. 8x - 10y = -100 50. -6x + 7y = 21 51. 12x + 15y = -45 52. -5x + 9y = –15 53. 16x + 11y = -88 54. 3x - 27y = 18

OpenStudy (anonymous):

49. -10y=-100-8x because y=mx+b try the next one

OpenStudy (anonymous):

whats the y and x intercept?

OpenStudy (anonymous):

and dont you have to solve for y?

OpenStudy (anonymous):

where x intercepts the graph, y is 0. so plug in y = 0 and solve for x. Then vice versa for y intercept

OpenStudy (anonymous):

but how do you 49?

OpenStudy (anonymous):

will it be y=10--8/10?

OpenStudy (anonymous):

its y=-8x/10+10

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

if your using the y=mx+b....you negate the 8x then you divide the -10 below the -8x and -100

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

m = y2 - y1/x2 - x1 (0, 10) and (12.5, 0)